Geometric Description

of the Connecting Homomorphism for Witt Groups

Paul Balmer^{1} and Baptiste Calmès^{2}

Received: December 10, 2008 Revised: October 23, 2009

Communicated by Ulf Rehmann

Abstract. We give a geometric setup in which the connecting homomorphism in the localization long exact sequence for Witt groups decomposes as the pull-back to the exceptional fiber of a suitable blow-up followed by a push-forward.

2000 Mathematics Subject Classification: Primary 19G12; Secondary 11E81. Keywords and Phrases: Witt group, localization sequence, connecting homomorphism, blow-up, push-forward, non-oriented theory.

Contents

1 Introduction  526
2 The regular case  529
3 The non-regular case  534
4 The main argument  537
5 The Main Theorem in the non-regular case  540
A Line bundles and dualizing complexes  543

1Research supported by SNSF grant PP002-112579 and NSF grant 0654397 2Research supported by EPSRC grant EP/E01786X/1

Documenta Mathematica 14 (2009) 525–550

1 Introduction

Witt groups form a very interesting cohomology theory in algebraic geometry. (For a survey, see [5].) Unlike the better known K-theory and Chow theory, Witt theory is not oriented in the sense of Levine-Morel [17] or Panin [22], as already visible on the non-standard projective bundle theorem, see Arason [2] and Walter [26]. Another way of expressing this is that push-forwards do not exist in sufficient generality for Witt groups. This “non-orientability” can make computations unexpectedly tricky. Indeed, the Witt groups of such elementary schemes as Grassmann varieties will appear for the first time in the companion article [6], whereas the corresponding computations for oriented cohomologies have been achieved more than 35 years ago in [16], using the well-known cellular decomposition of Grassmann varieties. See also [21] for general statements on cellular varieties. In oriented theories, there is a very useful computational technique, recalled in Theorem 1.3 below, which allows inductive computations for families of cellular varieties. Our paper originates in an attempt to extend this result to the nonoriented setting of Witt theory. Roughly speaking, such an extension is possible “half of the time”. In the remaining “half”, some specific ideas must come in and reflect the truly non-oriented behavior of Witt groups. To explain this rough statement, let us fix the setup, which will remain valid for the entire paper.

Setup 1.1. We denote by Sch the category of separated connected noetherian ℤ[\frac{1}{2}]-scheme. Let X, Z ∈ Sch be schemes and let ι : Z ↪ X be a regular closed immersion of codimension c ≥ 2. Let Bl = Bl_{Z}X be the blow-up of X along Z and E the exceptional fiber. Let U = X − Z ≅ Bl − E be the unaltered open complement. We have a commutative diagram

[figure: Z; ι; X; U; υ; υ̃; E; ι̃; π̃; Bl; π]  (1)

with the usual morphisms.

Consider now a cohomology theory with supports, say H^{∗}

⋯ ⟶^{∂} H_{Z}^{∗}(X) ⟶ H^{∗}(X) ⟶^{υ^{∗}} H^{∗}(U) ⟶^{∂} H_{Z}^{∗+1}(X) ⟶ ⋯  (2)

In this paper we shall focus on the case of Witt groups H^{∗} = W^{∗} but we take inspiration from H^{∗} being an oriented cohomology theory. Ideally, we would like conditions for the vanishing of the connecting homomorphism ∂ = 0 in the above localization long exact sequence. Even better would be conditions for the restriction υ^{∗} to be split surjective. When H^{∗} is an oriented theory, there is a well-known hypothesis under which such a splitting actually exists, namely :

Hypothesis 1.2. Assume that there exists an auxiliary morphism α̃ : Bl → Y

[figure: Z; ι; X; U; υ; υ̃; α; E; ι̃; π̃; Bl; π; α̃; Y]  (3)

such that α := α̃∘υ̃ : U → Y is an 𝔸^{∗}-bundle, i.e. every point of Y has a Zariski neighborhood over which α is isomorphic to a trivial 𝔸^{r}-bundle, for some r ≥ 0. See Ex. 1.5 for an explicit example with X, Y and Z being Grassmann varieties.

Theorem 1.3 (The oriented technique). Under Setup 1.1 and Hypothesis 1.2, assume X, Y and Z regular. Assume the cohomology theory H^{∗} is homotopy invariant for regular schemes and oriented, in that it admits push-forwards along proper morphisms satisfying flat base-change. Then, the restriction υ^{∗} : H^{∗}(X) → H^{∗}(U) is split surjective with explicit section π_{∗} ∘ α̃^{∗} ∘ (α^{∗})^{−1}, where π_{∗} : H^{∗}(Bl) → H^{∗}(X) is the push-forward. Hence the connecting homomorphism ∂ : H^{∗}(U) → H_{Z}^{∗+1}(X) vanishes and the above localization long exact sequence (2) reduces to split short exact sequences 0 → H_{Z}^{∗}(X) → H^{∗}(X) → H^{∗}(U) → 0.

Proof. By homotopy invariance, we have α^{∗} : H^{∗}(Y) →^{∼} H^{∗}(U). By basechange, υ^{∗}∘π_{∗} = υ̃^{∗} and since υ̃^{∗}∘α̃^{∗} = α^{∗}, we have υ^{∗}∘π_{∗}∘α̃^{∗}∘(α^{∗})^{−1} = id.

The dichotomy between the cases where the above technique extends to Witt groups and the cases where is does not, comes from the duality. To understand this, recall that one can consider Witt groups W^{∗}(X, L) with duality twisted by a line bundle L on the scheme X. Actually only the class of the twist L in Pic(X)/2 really matters since we have square-periodicity isomorphisms for all M ∈ Pic(X)

W^{∗}(X, L) ≅ W^{∗}(X, L ⊗ M^{⊗2}) .  (4)

Here is a condensed form of our Theorem 2.3 and Main Theorem 2.6 below :

Theorem 1.4. Under Hypothesis 1.2, assume X, Y and Z regular. Let L ∈ Pic(X). Then there exists an integer λ(L) ∈ ℤ (defined by (8) below) such that :

(A) If λ(L) ≡ c − 1 mod 2 then the restriction υ^{∗} : W^{∗}(X, L) → W^{∗}(U, L_{|U}) is split surjective with a section given by the composition π_{∗} ∘α̃^{∗} ∘(α^{∗})^{−1}. Hence the connecting homomorphism W^{∗}(U, L_{|U}) ⟶^{∂} W_{Z}^{∗+1}(X, L) vanishes and the localization long exact sequence reduces to split short exact sequences

0 ⟶ W_{Z}^{∗}(X, L) ⟶ W^{∗}(X, L) ⟶ W^{∗}(U, L_{|U}) ⟶ 0 .

(B) If λ(L) ≡ c mod 2 then the connecting homomorphism ∂ is equal to a composition of pull-backs and push-forwards : ∂ = ι_{∗} ∘π̃_{∗} ∘ι̃^{∗} ∘α̃^{∗} ∘(α^{∗})^{−1}.

This statement requires some explanations. First of all, note that we have used push-forwards for Witt groups, along π : Bl → X in (A) and along π̃ : E → Z and ι : Z → X in (B). To explain this, recall that the push-forward in Witt theory is only conditionally defined. Indeed, given a proper morphism f : X′ → X between (connected) regular schemes and given a line bundle L ∈ Pic(X), the push-forward homomorphism does not map W^{∗}(X′, f^{∗}L) into W^{∗}(X, L), as one could naively expect, but the second author and Hornbostel [8] showed that Grothendieck-Verdier duality yields a twist by the relative canonical line bundle ω_{f} ∈ Pic(X′) :

W^{i+dim(f)}( X′ , ω_{f} ⊗ f^{∗}L ) ⟶^{f_{∗}} W^{i}( X , L) .  (5)

Also note the shift by the relative dimension, dim(f) := dim X′ −dim X, which is not problematic, since we can always replace i ∈ ℤ by i − dim(f).

More trickily, if you are given a line bundle M ∈ Pic(X′) and if you need a push-forward W^{∗}(X′, M) → W^{∗−dim(f)}(X,?) along f : X′ → X, you first need to check that M is of the form ω_{f} ⊗ f^{∗}L for some L ∈ Pic(X), at least module squares. Otherwise, you simply do not know how to push-forward. This is precisely the source of the dichotomy of Theorem 1.4, as explained in Proposition 2.1 below.

At the end of the day, it is only possible to transpose to Witt groups the oriented technique of Theorem 1.3 when the push-forward π_{∗} exists for Witt groups. But actually, the remarkable part of Theorem 1.4 is Case (B), that is our Main Theorem 2.6 below, which gives a description of the connecting homomorphism ∂ when we cannot prove it zero by the oriented method. This is the part where the non-oriented behavior really appears. See more in Remark 2.7. Main Theorem 2.6 is especially striking since the original definition of the connecting homomorphism given in [3, § 4] does not have such a geometric flavor of pull-backs and push-forwards but rather involves abstract techniques of triangulated categories, like symmetric cones, and the like. Our new geometric description is also remarkably simple to use in applications, see [6]. Here is the example in question.

Example 1.5. Let k be a field of characteristic not 2. (We describe flag varieties over k by giving their k-points, as is customary.) Let 1 ≤ d ≤ n. Fix a codimension one subspace k^{n−1} of k^{n}. Let X = Gr_{d}(n) be the Grassmann variety of d-dimensional subspaces V_{d} ⊂ k^{n} and let Z ⊂ X be the closed subvariety of those subspaces V_{d} contained in k^{n−1}. The open complement U = X − Z consists of those V_{d} ⊄ k^{n−1}. For such V_{d} ∈ U, the subspace V_{d} ∩ k^{n−1} ⊂ k^{n−1} has dimension d − 1. This construction defines an 𝔸^{n−d}- bundle α : U → Y := Gr_{d−1}(n − 1), mapping V_{d} to V_{d} ∩ k^{n−1}. This situation relates the Grassmann variety X = Gr_{d}(n) to the smaller ones Z = Gr_{d}(n − 1) and Y = Gr_{d−1}(n − 1). Diagram (1) here becomes

\xymatrix{Gr_{d}(n − 1) \ar@{^{(}->}[r]^{ι} & Gr_{d}(n) & U \ar@{^{(}->}[l]_{υ} \ar@{^{(}->}[dl]_{υ̃} \ar[d]^{α} \\ E \ar@{^{(}->}[r]_{ι̃} \ar[u]^{π̃} & Bl \ar[u]^{π} \ar[r]_{α̃} & Gr_{d−1}(n − 1) .}

The blow-up Bl is the variety of pairs of subspaces V_{d−1} ⊂ V_{d} in k^{n}, such that V_{d−1} ⊂ k^{n−1}. The morphisms π : Bl → X and α̃ : Bl → Y forget V_{d−1} and V_{d} respectively. The morphism υ̃ maps V_{d} ⊄ k^{n−1} to the pair (V_{d} ∩ k^{n−1}) ⊂ V_{d}. Applying Theorem 1.3 to this situation, Laksov [16] computes the Chow groups of Grassmann varieties by induction. For Witt groups though, there are cases where the restriction W^{∗}(X, L) → W^{∗}(U, L_{|U}) is not surjective (see [6, Cor. 6.7]). Nevertheless, thank to our geometric description of the connecting homomorphism, we have obtained a complete description of the Witt groups of Grassmann varieties, for all shifts and all twists, to appear in [6]. In addition to the present techniques, our computations involve other ideas, specific to Grassmann varieties, like Schubert cells and desingularisations thereof, plus some combinatorial bookkeeping by means of special Young diagrams. Including all this here would misleadingly hide the simplicity and generality of the present paper. We therefore chose to publish the computation of the Witt groups of Grassmann varieties separately in [6].

The paper is organized as follows. Section 2 is dedicated to the detailed explanation of the above dichotomy and the proof of the above Case (A), see Theorem 2.3. We also explain Case (B) in our Main Theorem 2.6 but its proof is deferred to Section 5. The whole Section 2 is written, as above, under the assumption that all schemes are regular. This assumption simplifies the statements but can be removed at the price of introducing dualizing complexes and coherent Witt groups, which provide the natural framework over non-regular schemes. This generalization is the purpose of Section 3. There, we even drop the auxiliary Hypothesis 1.2, i.e. the dotted part of Diagram (3). Indeed, our Main Lemma 3.5 gives a very general description of the connecting homomorphism applied to a Witt class over U, if that class comes from the blow-up Bl via restriction υ̃^{∗}. The proof of Main Lemma 3.5 occupies Section 4. Finally, Hypothesis 1.2 re-enters the game in Section 5, where we prove our Main Theorem 2.6 as a corollary of a non-regular generalization given in Theorem 5.1. For the convenience of the reader, we gathered in Appendix A the needed results about Picard groups, canonical bundles and dualizing complexes, which are sometimes difficult to find in the literature. The conscientious reader might want to start with that appendix.

2 The regular case

We keep notation as in Setup 1.1 and we assume all schemes to be regular. This section can also be considered as an expanded introduction.

As explained after Theorem 1.4 above, we have to decide when the push-forward along π : Bl → X and along π̃ : E → Z exist. By (5), we need to determine the canonical line bundles ω_{π} ∈ Pic(Bl) and ωπ̃ ∈ Pic(E). This is classical and is recalled in Appendix A. First of all, Proposition A.6 gives

[figure: Pic; (; Z; ι; X; U; υ; υ̃; E; ι̃; π̃; Bl; π; ); ≅; Pic(Z); _{( \begin{matrix}1 \\ 0 \end{matrix} )}; Pic(X); ι^{∗}; _{( \begin{matrix}1 \\ 0 \end{matrix} )}; Pic(X) .; Pic(Z) ⊕ ℤ; Pic(X) ⊕ ℤ; _{( \begin{matrix}ι^{∗} & 0 \\ 0 & 1 \end{matrix} )}; (1 0)]

The ℤ summands in Pic(Bl) and Pic(E) are generated by O(E) = O_{Bl}(−1) and O(E)_{|E} = O_{E}(−1) respectively. Then Proposition A.11 gives the wanted

ω_{π} = (0, c − 1) in Pic(X) ⊕ ℤ ≅ Pic(Bl) and
ωπ̃ = (−ω_{ι}, c) in Pic(Z) ⊕ ℤ ≅ Pic(E) .

(6)

So, statistically, picking a line bundle M ∈ Pic(Bl) at random, there is a 50% chance of being able to push-forward W^{∗}(Bl, M) → W^{∗}(X, L) along π for some suitable line bundle L ∈ Pic(X). To justify this, observe that

coker ( [figure: Pic(X); π^{∗}; Pic(Bl)] )/2 ≅ ℤ/2

and tensoring by ω_{π} is a bijection, so half of the elements of Pic(Bl)/2 are of the form ω_{π} ⊗ π^{∗}(L). The same probability of 50% applies to the push forward along π̃ : E → Z but interestingly in complementary cases, as we summarize now.

Proposition 2.1. With the notation of 1.1, assume X and Z regular. Recall that c = codim_{X}(Z). Let M ∈ Pic(Bl). Let L ∈ Pic(X) and ℓ ∈ ℤ be such that M = (L, ℓ) in Pic(Bl) = Pic(X) ⊕ ℤ, that is, M = π^{∗}L ⊗ O(E)^{⊗ℓ}.

(A) If ℓ ≡ c − 1 mod 2, we can push-forward along π : Bl → X, as follows :

W^{∗}(Bl, M) ≅ W^{∗}(Bl, ω_{π} ⊗ π^{∗}L) →^{π_{∗}} W^{∗}(X, L) .

(B) If ℓ ≡ c mod 2, we can push-forward along π̃ : E → Z, as follows :

W^{∗}(E, M_{|E}) ≅ W^{∗}(E, ωπ̃ ⊗ π̃^{∗}(ω_{ι} ⊗ L_{|Z})) →^{π̃_{∗}} W^{∗−c+1}(Z, ω_{ι} ⊗ L_{|Z}).

In each case, the isomorphism ≅ comes from square-periodicity in the twist (4) and the subsequent homomorphism is the push-forward (5).

Proof. We only have to check the congruences in Pic /2. By (6), when ℓ ≡ c − 1 mod 2, we have [ω_{π} ⊗ π^{∗}L] = [(L, ℓ)] = [M] in Pic(Bl)/2. When ℓ ≡ c mod 2, we have [ωπ̃⊗π̃^{∗}(ω_{ι}⊗L_{|Z})] = [(L_{|Z} , ℓ)] = [M_{|E}] in Pic(E)/2. To apply (5), note that dim(π) = 0 since π is birational and dim(π̃) = c − 1 since E = ℙ_{Z}(C_{Z/X}) is the projective bundle of the rank-c conormal bundle C_{Z/X} over Z.

So far, we have only used Setup 1.1. Now add Hypothesis 1.2 with Y regular.

Remark 2.2. Since Picard groups of regular schemes are homotopy invariant, the 𝔸^{∗}-bundle α : U → Y yields an isomorphism α^{∗} : Pic(Y) →^{∼} Pic(U). Let us identify Pic(Y) with Pic(U), and hence with Pic(X) as we did above since c = codim_{X}(Z) ≥ 2. We also have O(E)_{|U} ≃ O_{U}. Putting all this together, the right-hand part of Diagram (3) yields the following on Picard groups :

[figure: Pic; (; X; U; υ; α; υ̃; Bl; π; α̃; Y; ); ≅; Pic(X); _{( \begin{matrix}1 \\ 0 \end{matrix} )}; Pic(X); Pic(X) ⊕ ℤ; (1 0); Pic(X) .; _{( \begin{matrix}1 \\ λ \end{matrix} )}]

Note that the lower right map Pic(X) ≅ Pic(Y) ⟶^{α̃^{∗}} Pic(Bl) ≅ Pic(X)⊕ℤ must be of the form ( \begin{matrix}1 \\ λ \end{matrix} ) by commutativity (i.e. since (^{1} ^{0} ) · ( \begin{matrix}1 \\ λ \end{matrix} ) = 1) but there is no reason for its second component λ : Pic(X) → ℤ to vanish. This is indeed a key observation. In other words, we have two homomorphisms from Pic(X) to Pic(Bl), the direct one π^{∗} and the circumvolant one α̃^{∗} ∘ (α^{∗})^{−1} ∘ υ^{∗} going via U and Y

[figure: Pic(X); υ^{∗}; ≃; π^{∗}; ≠; Pic(U); (α^{∗})^{−1}; ≃; Pic(Bl); Pic(Y); α̃^{∗}]  (7)

and they do not coincide in general. The difference is measured by λ, which depends on the choice of Y and on the choice of α̃ : Bl → Y, in Hypothesis 1.2. So, for every L ∈ Pic(X), the integer λ(L) ∈ ℤ is defined by the equation

α̃^{∗} (α^{∗})^{−1} υ^{∗}(L) = π^{∗}(L) ⊗ O(E)^{⊗λ(L)}  (8)

in Pic(Bl). Under the isomorphism Pic(Bl) ≅ Pic(X) ⊕ ℤ, the above equation can be reformulated as α̃^{∗} (α^{∗})^{−1} υ^{∗}(L) = (L, λ(L)).

Theorem 2.3 (Partial analogue of Theorem 1.3). With the notation of 1.1, assume Hypothesis 1.2 and assume X, Y, Z regular. Recall that c = codim_{X}(Z). Let L ∈ Pic(X) and consider the integer λ(L) ∈ ℤ defined in (8) above. If λ(L) ≡ c−1 mod 2 then the restriction υ^{∗} : W^{∗}(X, L) → W^{∗}(U, L_{|U}) is split surjective, with an explicit section given by the composition π_{∗} ∘ α̃^{∗} ∘ (α^{∗})^{−1}

\xymatrix{W^{∗}(X, L) &  & W^{∗}(U, L_{|U}) \ar[d]^{(α^{∗})^{−1}}_{≃} \\  & W^{∗}(Bl , ω_{π} ⊗ π^{∗}L) ≅ W^{∗}(Bl , α̃^{∗} (α^{∗})^{−1}L_{|U} ) \ar[ul]^{π_{∗}} & W^{∗}(Y , (α^{∗})^{−1} L_{|U} ) \ar[l]^{α̃^{∗}}}

Proof. The whole point is that π_{∗} can be applied after α̃^{∗} ∘ (α^{∗})^{−1}, that is, on W^{∗}(Bl , α̃^{∗} (α^{∗})^{−1} υ^{∗}(L)). This holds by Proposition 2.1 (A) applied to

M := α̃^{∗} (α^{∗})^{−1} υ^{∗}(L) =^{(8)} (L, λ(L)) ∈ Pic(X) ⊕ ℤ = Pic(Bl) .  (9)

The assumption λ(L) ≡ c − 1 mod 2 expresses the hypothesis of Proposition 2.1 (A). Checking that we indeed have a section goes as in the oriented case, see Thm. 1.3 :

υ^{∗} ∘ π_{∗} ∘ α̃^{∗} ∘ (α^{∗})^{−1} = υ̃^{∗} ∘ α̃^{∗} ∘ (α^{∗})^{−1} = α^{∗} ∘ (α^{∗})^{−1} = id .

The first equality uses base-change [8, Thm. 6.9] on the left-hand cartesian square :

\xymatrix{X & U \ar@{^{(}->}[l]_{υ} \\ Bl \ar[u]^{π} & U \ar@{^{(}->}[l]_{υ̃} \ar[u]_{id}}  L  L_{|U} ω_{π} ⊗ π^{∗}L  L_{|U}

with respect to the right-hand line bundles. Note that (ω_{π})_{|U} = O_{U} by (6).

Remark 2.4. In the above proof, see (9), we do not apply Proposition 2.1 to M being π^{∗}L, as one could first expect; see Remark 2.2. Consequently, our condition on L, namely λ(L) ≡ c − 1 mod 2, does not only depend on the codimension c of Z in X but also involves (hidden in the definition of λ) the particular choice of the auxiliary scheme Y and of the morphism α̃ : Bl → Y of Hypothesis 1.2.

Remark 2.5. The legitimate question is now to decide what to do in the remaining case, that is, when λ(L) ≡ c mod 2. As announced, this is the central goal of our paper (Thm. 2.6 below). So, let L ∈ Pic(X) be a twist such that push-forward along π : Bl → X cannot be applied to define a section to the restriction W^{∗}(X, L) → W^{∗}(U, L_{|U}) as above. Actually, we can find examples of such line bundles for which this restriction is simply not surjective (see Ex. 1.5). The natural problem then becomes to compute the possibly nonzero connecting homomorphism ∂ : W^{∗}(U, L_{|U}) → W_{Z}^{∗+1}(X, L). Although not absolutely necessary, it actually simplifies the formulation of Theorem 2.6 to use dévissage from [9, §6], i.e. the fact that push-forward along a regular closed immersion is an isomorphism

ι_{∗} : W^{∗−c}(Z, ω_{ι} ⊗ L_{|Z}) →^{∼} W_{Z}^{∗}(X, L) .  (10)

Using this isomorphism, we can replace the Witt groups with supports by Witt groups of Z in the localization long exact sequence, and obtain a long exact sequence

\xymatrix{⋯ & W^{∗}(X, L) \ar[r]^{υ^{∗}} & W^{∗}(U, L_{|U}) \ar@{-->}[ddr]_{∂} \ar[r]^{∂′} & W^{∗+1−c}(Z, ω_{ι} ⊗ L_{|Z}) \ar[r] \ar[dd]_{≃}^{ι_{∗}} & W^{∗+1}(X, L) & ⋯ \\ (11) \\  &  &  & W_{Z}^{∗+1}(X, L) \ar[uur]}

We now want to describe ∂′ when λ(L) ≡ c mod 2 (otherwise ∂′ = 0 by Thm. 2.3). By the complete dichotomy of Proposition 2.1, we know that when the push-forward π_{∗} : W^{∗}(Bl, M) → W^{∗}(X,?) does not exist, here for M = α̃^{∗} (α^{∗})^{−1} υ^{∗}(L) by (9), then the following composition π̃_{∗} ∘ ι̃^{∗} exists and starts from the very group where π_{∗} cannot be defined and arrives in the very group where ∂′ itself arrives :

W^{∗}(Bl, M) ⟶^{ι̃^{∗}} W^{∗}(E, M_{|E}) ⟶^{π̃_{∗}} W^{∗−c+1}(Z, ω_{ι} ⊗ L_{|Z}) .

Hence, in a moment of exaltation, if we blindly apply this observation at the precise point where the oriented technique fails for Witt groups, we see that when we cannot define a section to restriction by the formula π_{∗} ∘(α̃^{∗} ∘(α^{∗})^{−1}) we can instead define a mysterious homomorphism (π̃_{∗} ∘ ι̃^{∗}) ∘ (α̃^{∗} ∘ (α^{∗})^{−1}).

Theorem 2.6 (Main Theorem in regular case). With the notation of 1.1, assume Hypothesis 1.2 and assume X, Y, Z regular. Let L ∈ Pic(X) and recall the integer λ(L) ∈ ℤ defined by (8). If λ(L) ≡ c mod 2 then the composition π̃_{∗} ∘ ι̃^{∗} ∘ α̃^{∗} ∘ (α^{∗})^{−1} is equal to the connecting homomorphism ∂′ of (11), that is, the following diagram commutes :

\xymatrix{W^{∗+1−c}(Z, ω_{ι} ⊗ L_{|Z}) &  & W^{∗}(U, L_{|U}) \ar[d]^{(α^{∗})^{−1}}_{≃} \ar[ll]_{∂′} \\  & W^{∗}(E , ωπ̃ ⊗ π̃^{∗}(ω_{ι} ⊗ L_{|Z}))≅W^{∗}(E , ι̃^{∗} α̃^{∗} (α^{∗})^{−1}L_{|U} ) \ar[ul]^{π̃_{∗}} & W^{∗}(Y, (α^{∗})^{−1} L_{|U} ) \ar[l]^{_{_{ι̃^{∗}} _{α̃^{∗}}}}}

This statement implies Thm. 1.4 (B) since ∂ = ι_{∗} ∂′ by (11). Its proof will be given after generalization to the non-regular setting, at the end of Section 5.

Remark 2.7. Let us stress the peculiar combination of Theorem 2.3 and Theorem 2.6. Start with a Witt class w_{U} over the open U ⊂ X, for the duality twisted by some L ∈ Pic(U) = Pic(X), and try to extend w_{U} to a Witt class w_{X} over X:

[figure: ∂′(w_{U}); ←50%|50%→; w_{X}; υ^{∗}; w_{U}; (α^{∗})^{−1}; w_{E}; π̃_{∗}; w_{Bl}; ι̃^{∗}; π_{∗}; w_{Y}; α̃^{∗}]

Then, either we can apply the same construction as for oriented theories, i.e. push-forward the class w_{Bl} := α̃^{∗} ∘ (α^{∗})^{−1}(w_{U}) from Bl to X along π, constructing in this way an extension w_{X} := π_{∗}(w_{Bl}) as wanted, or this last push-forward π_{∗} is forbidden on w_{Bl} because of the twist, in which case the Witt class w_{U} might simply not belong to the image of restriction υ^{∗}. The latter means that w_{U} might have a non-zero boundary ∂′(w_{U}) over Z, which then deserves to be computed. The little miracle precisely is that in order to compute this ∂′(w_{U}), it suffices to resume the above process where it failed, i.e. with w_{Bl}, and, since we cannot push it forward along π, we can consider the bifurcation of Proposition 2.1 and restrict this class w_{Bl} to the exceptional fiber E, say w_{E} := ι̃^{∗}w_{Bl}, and then push it forward along π̃. Of course, this does not construct an extension of w_{U} anymore, since this new class π̃_{∗}(w_{E}) lives over Z, not over X. Indeed, there is no reason a priori for this new class to give anything sensible at all. Our Main Theorem is that this construction in fact gives a formula for the boundary ∂′(w_{U}). Bottom line : Essentially the same geometric recipe of pull-back and pushforward either splits the restriction or constructs the connecting homomorphism. In particular, the connecting homomorphism is explicitly described in both cases.

3 The non-regular case

In Section 2, we restricted our attention to the regular case in order to grasp the main ideas. However, most results can be stated in the greater generality of separated and noetherian ℤ[\frac{1}{2}]-schemes admitting a dualizing complex. The goal of this section is to provide the relevant background and to extend Theorem 1.4 to this non-regular setting, see Main Lemma 3.5.

Remark 3.1. The coherent Witt groups W̃^{∗}(X, K_{X}) of a scheme X ∈ Sch (see 1.1) are defined using the derived category D_{coh}^{b}(X) of complexes of O_{X}- modules whose cohomology is coherent and bounded. Since X is noetherian and separated, D_{coh}^{b}(X) is equivalent to its subcategory D^{b}(coh(X)) of bounded complexes of coherent O_{X}-modules; see for instance [8, Prop. A.4]. The duality is defined using the derived functor RHom(−, K_{X}) where K_{X} ∈ D_{coh}^{b}(X) is a dualizing complex (see [19, § 3] or [8, § 2]), meaning that the functor RHom(−, K_{X}) defines a duality on D_{coh}^{b}(X). For example, a scheme is Gorenstein if and only if O_{X} itself is an injectively bounded dualizing complex and, in that case, all other dualizing complexes are shifted line bundles (see Lemma A.7). Regular schemes are Gorenstein, and for them, coherent Witt groups coincide with the usual “locally free” Witt groups W^{∗}(X, L) (i.e. the ones defined using bounded complexes of locally free sheaves instead of coherent ones). For any line bundle L, we still have a square-periodicity isomorphism

W̃(X, K_{X}) ≅ W̃(X, K_{X} ⊗ L^{⊗2})  (12)

given by the multiplication by the class in W^{0}(X, L^{⊗2}) of the canonical form L → L^{∨}⊗L^{⊗2}, using the pairing between locally free and coherent Witt groups. For any closed embedding Z ↪ X with open complement υ : U ↪ X, the restriction K_{U} := υ^{∗}K_{X} is a dualizing complex [19, Thm. 3.12] and the general triangulated framework of [3] gives a localization long exact sequence

⋯ ⟶^{∂} W̃_{Z}^{∗}(X, K_{X}) ⟶ W̃^{∗}(X, K_{X}) ⟶ W̃^{∗}(U, K_{U}) ⟶^{∂} W̃_{Z}^{∗+1}(X, K_{X}) ⟶ ⋯

(13)

As for K-theory, no such sequence holds in general for singular schemes and locally free Witt groups.

Remark 3.2. For coherent Witt groups, the push-forward along a proper morphism f : X′ → X takes the following very round form : If K_{X} is a dualizing complex on X then f^{!}K_{X} is a dualizing complex on X′ ([8, Prop. 3.9]) and the functor R f_{∗} : D_{coh}^{b}(X′) → D_{coh}^{b}(X) induces a push-forward ([8, Thm. 4.4])

f_{∗} : W̃^{i}(X′, f^{!}K_{X}) → W̃^{i}(X, K_{X}) .  (14)

Recall that f^{!} : D_{Qcoh}(X) → D_{Qcoh}(X′) is the right adjoint of R f_{∗}. If we twist the chosen dualizing complex K_{X} by a line bundle L ∈ Pic(X), this is transported to X′ via the following formula (see [8, Thm. 3.7])

f^{!}(K_{X} ⊗ L) ≃ f^{!}(K_{X}) ⊗ f^{∗}L .  (15)

In the regular case, push-forward maps are also described in Nenashev [20].

Remark 3.3. Let us also recall from [8, Thm. 4.1] that the pull-back

f^{∗} : W̃^{i}(X, K_{X}) → W̃^{i}(X′, Lf^{∗}K_{X})

along a finite Tor-dimension morphism f : X′ → X is defined if Lf^{∗}(K_{X}) is a dualizing complex (this is not automatically true). Together with the push-forward, this pull-back satisfies the usual flat base-change formula (see [8, Thm. 5.5]). A regular immersion f : X′ ↪ X has finite Tor-dimension since it is even perfect (see [1, p. 250]). Moreover, in that case, Lf^{∗} is the same as f^{!} up to a twist and a shift (see Proposition A.8), hence it preserves dualizing complexes.

Proposition 3.4. In Setup 1.1, let K_{X} be a dualizing complex on X. Let L ∈ Pic(X) and ℓ ∈ ℤ. Then K = π^{!}(K_{X}) ⊗ π^{∗}L ⊗ O(E)^{⊗ℓ} is a dualizing complex on Bl and any dualizing complex has this form, for some L ∈ Pic(X) and ℓ ∈ ℤ. Moreover, the dichotomy of Proposition 2.1 here becomes :

(A) If ℓ ≡ 0 mod 2, we can push-forward along π : Bl → X, as follows :

W̃^{∗}(Bl, K) ≅ W̃^{∗}(Bl, π^{!}(K_{X} ⊗ L)) →^{π_{∗}} W̃^{∗}(X, K_{X} ⊗ L) .

(B) If ℓ ≡ 1 mod 2, we can push-forward along π̃ : E → Z, as follows :

W̃^{∗}(E, Lι̃^{∗}K) ≅ W̃^{∗+1}(E, π̃^{!}ι^{!}(K_{X} ⊗ L))) →^{π̃_{∗}} W̃^{∗+1}(Z, ι^{!}(K_{X} ⊗ L)).

As before, in both cases, the first isomorphism ≅ comes from squareperiodicity (12) and the second morphism is push-forward (14).

Proof. The complex K_{Bl} := π^{!}K_{X} is a dualizing complex on Bl by Remark 3.2. By Lemma A.7 and Proposition A.6 (i), all dualizing complexes on Bl are of the form K = π^{!}(K_{X}) ⊗ π^{∗}L ⊗ O(E)^{⊗ℓ}, for unique L ∈ Pic(X) and ℓ ∈ ℤ. We only need to check the relevant parity for applying (12). Case (A) follows easily from (15) by definition of K and parity of ℓ. In (B), we need to compare Lι̃^{∗}K and π̃^{!}ι^{!}(K_{X} ⊗ L)[1]. By Proposition A.11 (iv), we know that ι̃^{!}(−) ≅ ι̃^{∗}O(E)[−1] ⊗ Lι̃^{∗}(−). We apply this and (15) in the second equality below, the first one using simply that ιπ̃ = πι̃:

π̃^{!}ι^{!}(K_{X} ⊗ L)[1] ≅ ι̃^{!}π^{!}(K_{X} ⊗ L)[1] ≅ ι̃^{∗}O(E)[−1] ⊗ Lι̃^{∗}(π^{!}(K_{X}) ⊗ π^{∗}L)[1] ≅
≅ ι̃^{∗}O(E) ⊗ Lι̃^{∗}(K ⊗ O(E)^{⊗−ℓ}) ≅ ι̃^{∗}O(E)^{⊗(1−ℓ)} ⊗ Lι̃^{∗}K.

Since 1 − ℓ is even, ι̃^{∗}O(E)^{⊗(1−ℓ)} is a square, as desired.

We now want to give the key technical result of the paper, which is an analogue of Theorem 1.4 in the non-regular setting. The idea is to describe the connecting homomorphism on Witt classes over U which admit an extension to the blow-up Bl. The key fact is the existence of an additional twist on Bl, namely the twist by O(E), which disappears on U (see A.1) and hence allows Case (B) below.

Main Lemma 3.5. In Setup 1.1, assume that X has a dualizing complex K_{X} and let K_{U} = υ^{∗}(K_{X}) and K_{Bl} = π^{!}(K_{X}); see Remarks 3.1 and 3.2. Let i ∈ ℤ.

(A) The following composition vanishes :

\xymatrix{W̃^{i}(Bl, K_{Bl}) \ar[r]^{υ̃^{∗}} & W̃^{i}(U, K_{U}) \ar[r]^{∂} & W̃_{Z}^{i+1}(X, K_{X}) .}

(B) The following composition (well-defined since υ̃^{∗}O(E) ≃ O_{U})

\xymatrix{W̃^{i}(Bl, K_{Bl} ⊗ O(E)) \ar[r]^{υ̃^{∗}} & W̃^{i}(U, K_{U} ⊗ υ̃^{∗}O(E))≅W̃^{i}(U, K_{U}) \ar[r]^{∂} & W̃_{Z}^{i+1}(X, K_{X})}

coincides with the composition

[figure: W̃^{i}(Bl, O(E) ⊗ K_{Bl}); ι̃^{∗}; W̃_{Z}^{i+1}(X, K_{X}); W̃^{i}(E, Lι̃^{∗}(O(E) ⊗ K_{Bl})); ≅; W̃^{i+1}(E, π̃^{!}ι^{!}K_{X}); π̃_{∗}; W̃^{i+1}(Z, ι^{!}K_{X}); ι_{∗}]

where the latter isomorphism ≅ is induced by the composition

Lι̃^{∗}(O(E) ⊗ K_{Bl}) ≅ ι̃^{∗}(O(E)) ⊗ Lι̃^{∗}(K_{Bl}) ≅ ι̃^{!}K_{Bl}[1] ≅ π̃^{!}ι^{!}K_{X}[1] . (16)

The proof of this result occupies Section 4. Here are just a couple of comments on the statement. Let us first of all explain the announced sequence of isomorphisms (16). The first one holds since Lι̃^{∗} is a tensor functor and since O(E) is a line bundle (hence is flat). The second one holds by Proposition A.9 (v). The last one follows by definition of K_{Bl} and the fact that ι π̃ = π ι̃. Finally, note that we use the pull-back ι̃^{∗} on coherent Witt groups as recalled in Remark 3.3.

4 The main argument

Surprisingly enough for a problem involving the blow-up Bl = Bl_{Z}(X) of X along Z, see (1), the case where codim_{X}(Z) = 1 is also interesting, even though, of course, in that case Bl = X and E = Z. In fact, this case is crucial for the proof of Main Lemma 3.5 and this is why we deal with it first. In the “general” proof where codim_{X}(Z) is arbitrary, we will apply the case of codimension one to ι̃: E ↪ Bl. Therefore, we use the following notation to discuss codimension one.

Notation 4.1. Let B ∈ Sch be a scheme with a dualizing complex K_{B} and ι̃: E ↪ B be a prime divisor, that is, a regular closed immersion of codimension one, of a subscheme E ∈ Sch. Let O(E) be the line bundle on B associated to E (see Definition A.1). Let υ̃ : U ↪ B be the open immersion of the open complement

\xymatrix{E \ar@{^{(}->}[r]^{ι̃} & B & U \ar@{^{(}->}[l]_{υ̃}}

U = B − E and let K_{U} be the dualizing complex υ̃^{∗}(K_{B}) on U.

Lemma 4.2 (Main Lemma in codimension one). With Notation 4.1, let i ∈ ℤ. Then :

(A) The composition

\xymatrix{W̃^{i}(B, K_{B}) \ar[r]^{υ̃^{∗}} & W̃^{i}(U, K_{U}) \ar[r]^{∂} & W̃_{E}^{i+1}(B, K_{B})}

is zero.

(B) The composition

\xymatrix{W̃^{i}(B, K_{B} ⊗ O(E)) \ar[r]^{υ̃^{∗}} & W̃^{i}(U, K_{U} ⊗ υ̃^{∗}O(E))≅W̃^{i}(U, K_{U}) \ar[r]^{∂} & W̃_{E}^{i+1}(B, K_{B})}

coincides with the composition

[figure: W̃^{i}(B, O(E) ⊗ K_{B}); ι̃^{∗}; W̃_{E}^{i+1}(B, K_{B}); W̃^{i}(E, Lι̃^{∗}(O(E) ⊗ K_{B})); ≅; W̃^{i}(E, ι̃^{!}K_{B}[1]); ≅; W̃^{i+1}(E, ι̃^{!}K_{B}); ι̃_{∗}]

where the first isomorphism ≅ is induced by the following isomorphism

Lι̃^{∗}(O(E) ⊗ K_{B}) ≅ ι̃^{∗}(O(E)) ⊗ Lι̃^{∗}(K_{B}) ≅ ι̃^{!}(K_{B})[1].  (17)

Proof. Case (A) is simple : The composition of two consecutive morphisms in the localization long exact sequence (13) is zero. Case (B) is the nontrivial one. The isomorphisms (17) are the same as in (16).

At this stage, we upload the definition of the connecting homomorphism for Witt groups ∂ : W̃^{i}(U, K_{U}) → W̃_{E}^{i+1}(B, K_{B}), which goes as follows : Take a non-degenerate symmetric space (P, ϕ) over U for the i^{th}-shifted duality with values in K_{U} ; there exists a possibly degenerate symmetric pair (Q, ψ) over B for the same duality (with values in K_{B}) which restricts to (P, ϕ) over U ; compute its symmetric cone d(Q, ψ), which is essentially the cone of ψ equipped with a natural metabolic form; see [3, § 4] or [4, Def. 2.3] for instance ; for any choice of such a pair (Q, ψ), the boundary ∂(P, ϕ) ∈ W̃_{E}^{i+1}(B, K_{B}) is the Witt class of d(Q, ψ). There is nothing really specific to dualizing complexes here. The above construction is a purely triangular one, as long as one uses the same duality for the ambient scheme B, for the open U ⊂ B and for the Witt group of B with supports in the closed complement E. The subtlety of statement (B) is that we start with a twisted duality on the scheme B which is not the duality used for ∂, but which agrees with it on U by the first isomorphism ≅ in statement (B). Now, take an element in W̃^{i}(B, O(E) ⊗ K_{B}). It is the Witt-equivalence class of a symmetric space (P, ϕ) over B with respect to the i^{th}-shifted duality with values in O(E) ⊗ K_{B}. The claim of the statement is that, modulo the above identifications of dualizing complexes, we should have

∂(υ̃^{∗}(P, ϕ)) = ι̃_{∗}(ι̃^{∗}(P, ϕ))  (18)

in W̃_{E}^{i+1}(B, K_{B}). By the above discussion, in order to compute ∂(υ̃^{∗}(P, ϕ)), we need to find a symmetric pair (Q, ψ) over B, for the duality given by K_{B}, and such that υ̃^{∗}(Q, ψ) = υ̃^{∗}(P, ϕ). Note that we cannot take for (Q, ψ) the pair (P, ϕ) itself because (P, ϕ) is symmetric for the twisted duality O(E)⊗K_{B} on B. Nevertheless, it is easy to “correct” (P, ϕ) as follows. As in Definition A.1, we have a canonical homomorphism of line bundles :

σ_{E} : O(E)^{∨} → O_{B} .

The pair (O(E)^{∨}, σ_{E}) is symmetric in the derived category D^{b}(VB(B)) of vector bundles over B, with respect to the 0^{th}-shifted duality twisted by O(E)^{∨}, because the target of σ_{E} is the dual of its source : (O(E)^{∨})^{∨}[0] ⊗ O(E)^{∨} ≅ O_{B}. Let us define the wanted symmetric pair (Q, ψ) in D_{coh}^{b}(B) for the i^{th}-shifted duality with values in K_{B} as the following product :

(Q, ψ) := (O(E)^{∨}, σ_{E}) ⊗ (P, ϕ) .

Note that we tensor a complex of vector bundles with a coherent one to get a coherent one, following the formalism of [4, § 4] where such external products are denoted by ⋆. We claim that the restriction of (Q, ψ) to U is nothing but υ̃^{∗}(P, ϕ). This is easy to check since O(E)_{|U} = O_{U} via σ_{E} (see A.1), which means (O(E)^{∨}, σ_{E})_{|U} = 1_{U}. So, by the construction of the connecting homomorphism ∂ recalled at the beginning of the proof, we know that ∂(υ̃^{∗}(P, ϕ)) can be computed as d(Q, ψ). This reads :

∂(υ̃^{∗}(P, ϕ)) = d((O(E)^{∨}, σ_{E}) ⊗ (P, ϕ)).

Now, we use that (P, ϕ) is non-degenerate and that therefore (see [4, Rem. 5.4] if necessary) we can take (P, ϕ) out of the above symmetric cone d(...), i.e.

∂(υ̃^{∗}(P, ϕ)) = d(O(E)^{∨}, σ_{E}) ⊗ (P, ϕ) .  (19)

Let us compute the symmetric cone d(O(E)^{∨}, σ_{E}) =: (C, χ). Note that this only involves vector bundles. We define C to be the cone of σ_{E} and we equip it with a symmetric form χ : C →^{∼} C^{∨}[1] ⊗ O(E)^{∨} for the duality used for (O(E), σ_{E}) but shifted by one, that is, for the 1^{st} shifted duality with values in O(E)^{∨}. One checks that (C, χ) is given by the following explicit formula :

\xymatrix{C = \ar[d]_{χ} & (⋯ \ar[r] & 0 \ar[r] \ar[d] & 0 \ar[r] \ar[d] & O(E)^{∨} \ar[r]^{σ_{E}} \ar[d]_{−1} & O_{B} \ar[r] \ar[d]^{1} & 0 \ar[r] \ar[d] & 0 \ar[r] \ar[d] & ⋯) \\ C^{∨}[1] ⊗ O(E)^{∨} = & _{(⋯} \ar[r] & 0 \ar[r] & 0 \ar[r] & O(E)^{∨} \ar[r]_{−(σ_{E})^{∨}} & O_{B} \ar[r] & 0 \ar[r] & 0 \ar[r] & _{⋯)}}

(20)

where the complexes have O_{B} in degree zero. Now, observe that the complex C is a resolution of ι̃_{∗}(O_{E}) over B, by Definition A.1, that is, C ≃ ι̃_{∗}(O_{E}) in the derived category of B. Moreover, by Propositions A.8 and A.9 (ii), we have ι̃^{!}(O(E)^{∨}[1]) = ωι̃[−1] ⊗ ι̃^{∗}(O(E)^{∨}[1]) ≅ O_{E}. Using this, one checks the conceptually obvious fact that χ is also the push-forward along the perfect morphism ι̃ of the unit form on O_{E}. See Remark 4.3 below for more details. This means that we have an isometry in D_{E}^{b}(VB(B))

d(O(E)^{∨}, σ_{E}) = ι̃_{∗}(1_{E})

of symmetric spaces with respect to the 1^{st} shifted duality with values in O(E)^{∨}. Plugging this last equality in (19), and using the projection formula (see Remark 4.3) we obtain

∂(υ̃^{∗}(P, ϕ)) = ι̃_{∗}(1_{E}) ⊗ (P, ϕ) = ι̃_{∗}(1_{E} ⊗ ι̃^{∗}(P, ϕ)) = ι̃_{∗}(ι̃^{∗}(P, ϕ)) .

This is the claimed equality (18).

Remark 4.3. In the above proof, we use the “conceptually obvious fact” that the push-forward of the unit form on O_{E} is indeed the χ of (20). This fact is obvious to the expert but we cannot provide a direct reference for this exact statement. However, if the reader does not want to do this lengthy verification directly, the computation of [8, § 7.2] can be applied essentially verbatim. The main difference is that here, we are considering a push-forward of locally free instead of coherent Witt groups along a regular embedding. Such a pushforward is constructed using the same tensor formalism as the proper pushforwards for coherent Witt groups considered in loc. cit. along morphisms that are proper, perfect and Gorenstein, which is true of a regular embedding. In loc. cit. there is an assumption that the schemes are Gorenstein, ensuring that the line bundles are dualizing complexes. But here, the dualizing objects for our category of complexes of locally free sheaves are line bundles anyway and the extra Gorenstein assumption is irrelevant. Moreover, the projection formula used in the above proof is established in complete generality for non necessarily regular schemes by the same method as in [8, § 5.7] using the pairing between the locally free derived category and the coherent one to the coherent one. More precisely, this pairing is just a restriction of the quasi-coherent pairing D_{Qcoh} × D_{Qcoh} ⟶^{⊗} D_{Qcoh} of loc. cit. to these subcategories. By the general tensor formalism of [10], for any morphism f : X → Y as above, for any object A (resp. B) in the quasi-coherent derived category of X (resp. Y), we obtain a projection morphism in D_{Qcoh}(Y)

R f_{∗}(A) ⊗ B ⟶ R f_{∗}(A ⊗ Lf^{∗}(B)),

see [10, Prop. 4.2.5]. It is an isomorphism by [8, Thm. 3.7]. We actually only use it for A a complex of locally free sheaves and B a complex with coherent and bounded cohomology. The projection formula is implied by [10, Thm. 5.5.1].

Proof of Main Lemma 3.5. Case (A) follows from the codimension one case and the compatibility of push-forwards with connecting homomorphisms (here along the identity of U). Case (B) follows from the outer commutativity of the following diagram :

[figure: W̃^{i}(U, K_{U}); ∂; W̃_{Z}^{i+1}(X, K_{X}); W̃^{i+1}(Z, ι^{!}K_{X}); ι_{∗}; W̃^{i}(U, K_{U}); ∂; W̃_{E}^{i+1}(Bl, K_{Bl}); π_{∗}; W̃^{i+1}(E, π̃^{!}ι^{!}K_{X}); ι̃_{∗}; π̃_{∗}; W̃^{i}(Bl, O(E) ⊗ K_{Bl}); υ̃^{∗}; ι̃^{∗}; W̃^{i}(Bl, Lι̃^{∗}(O(E) ⊗ K_{Bl})); ≅; W̃^{i+1}(E, ι̃^{!}K_{Bl}); ≅]

(21)

We shall now verify the inner commutativity of this diagram. The upper left square of (21) commutes by compatibility of push-forward with connecting homomorphisms. The upper right square of (21) simply commutes by functoriality of push-forward applied to ι ∘ π̃ = π ∘ ι̃. Most interestingly, the lower part of (21) commutes by Lemma 4.2 applied to the codimension one inclusion ι̃: E ↪ Bl.

5 The Main Theorem in the non-regular case

Without regularity assumptions, we have shown in Main Lemma 3.5 how to compute the connecting homomorphism ∂ : W̃^{∗}(U, K_{U}) → W̃_{Z}^{∗+1}(X, K_{X}) on those Witt classes over U which come from Bl = Bl_{Z}(X) by restriction υ̃^{∗}. The whole point of adding Hypothesis 1.2 is precisely to split υ̃^{∗}, that is, to construct for each Witt class on U an extension on Bl. In the regular case, this follows from homotopy invariance of Picard groups and Witt groups. In the non-regular setting, things are a little more complicated. Let us give the statement and comment on the hypotheses afterwards (see Remark 5.2).

Main Theorem 5.1. In Setup 1.1, assume that X has a dualizing complex K_{X} and equip U with the restricted complex K_{U} = υ^{∗}(K_{X}). Assume Hypothesis 1.2 and further make the following hypotheses :

(a) There exists a dualizing complex K_{Y} on Y such that α^{∗}K_{Y} = K_{U}.

(b) The 𝔸^{∗}-bundle α induces an isomorphism W̃^{∗}(Y, K_{Y}) →^{∼} W̃^{∗}(U, K_{U}).

(c) The morphism α̃ is of finite Tor dimension and Lα̃^{∗}(K_{Y}) is dualizing.

(d) Sequence (25) is exact : ℤ → Pic(Bl) → Pic(U). (See Proposition A.3.)

Then Lα̃^{∗}(K_{Y}) ≃ π^{!}K_{X} ⊗ O(E)^{⊗n} for some n ∈ ℤ, and the following holds true :

(A) If n can be chosen even, the composition π_{∗}α̃^{∗}(α^{∗})^{−1} is a section of υ^{∗}.

(B) If n can be chosen odd, the composition ι_{∗}π̃_{∗}ι̃^{∗}α̃^{∗}(α^{∗})^{−1} coincides with the connecting homomorphism ∂ : W̃^{∗}(U, K_{U}) → W̃_{E}^{∗+1}(X, K_{X}).

Proof. By (c) and Remark 3.2 respectively, both Lα̃^{∗}(K_{Y}) and π^{!}K_{X} are dualizing complexes on Bl. By Lemma A.7 (i), they differ by a shifted line bundle : Lα̃^{∗}(K_{Y}) ≃ π^{!}K_{X} ⊗ L[m] with L ∈ Pic(Bl) and m ∈ ℤ. Restricting to U, we get

K_{U} ⊗ υ̃^{∗}L[m] ≃ υ̃^{∗}π^{!}K_{X} ⊗ υ̃^{∗}L[m]

≃ υ̃^{∗}(π^{!}K_{X} ⊗ L[m]) ≃ υ̃^{∗} Lα̃^{∗}(K_{Y}) ≃ α^{∗}K_{Y} ≃ K_{U}

where the first equality holds by flat base-change ([8, Thm. 5.5]). Thus, υ̃^{∗}L[m] is the trivial line bundle on U by Lemma A.7 (ii). So m = 0 and, by (d), L ≃ O(E)^{⊗n} for some n ∈ ℤ. This gives Lα̃^{∗}(K_{Y}) ≃ π^{!}K_{X} ⊗ O(E)^{⊗n} as claimed. We now consider coherent Witt groups. By (c) and Remark 3.3, α̃ induces a morphism α̃^{∗} : W̃^{∗}(Y, K_{Y}) → W̃(Bl, Lα̃^{∗}K_{Y}). By Lemma A.12, the flat morphism α induces a homomorphism α^{∗} : W̃^{∗}(Y, K_{Y}) → W̃^{∗}(U, K_{U}) which is assumed to be an isomorphism in (b). So, we can use (α^{∗})^{−1}. When n is even, we have

υ^{∗}π_{∗}α̃^{∗}(α^{∗})^{−1} = υ̃^{∗}α̃^{∗}(α^{∗})^{−1} = α^{∗}(α^{∗})^{−1} = id

where the first equality holds by flat base-change ([8, Thm. 5.5]). This proves (A). On the other hand, when n is odd, we have

ι_{∗}π̃_{∗}ι̃^{∗}α̃^{∗}(α^{∗})^{−1} = ∂ υ̃^{∗}α̃^{∗}(α^{∗})^{−1} = ∂ α^{∗}(α^{∗})^{−1} = ∂

where the first equality holds by Main Lemma 3.5 (B).

Remark 5.2. Hypothesis (a) in Theorem 5.1 is always true when Y admits a dualizing complex and homotopy invariance holds over Y for the Picard group (e.g. Y regular). Homotopy invariance for coherent Witt groups should hold in general but only appears in the literature when Y is Gorenstein, see Gille [12]. This means that Hypothesis (b) is a mild one. Hypothesis (d) is discussed in Proposition A.3.

Remark 5.3. In Theorem 5.1, the equation Lα̃^{∗}(K_{Y}) ≃ π^{!}K_{X} ⊗ O(E)^{⊗n}, for n ∈ ℤ, should be considered as a non-regular analogue of Equation (8). In Remark 2.2, we discussed the compatibility of the various lines bundles on the schemes X, U, Y and Bl. Here, we need to control the relationship between dualizing complexes instead and we do so by restricting to U and by using the exact sequence (25). Alternatively, one can remove Hypothesis (d) and directly assume the relation Lα̃^{∗}(K_{Y}) ≃ π^{!}(K_{X})⊗O(E)^{⊗n} for some n ∈ ℤ. This might hold in some particular examples even if (25) is not exact.

For the convenience of the reader, we include the proofs of the following facts.

Lemma 5.4. If X is Gorenstein, then Z and Bl are Gorenstein. If X is regular, Bl is regular.

Proof. By Prop. A.8, π^{!}(O_{X}) is the line bundle ω_{π}. Since π is proper, π^{!} preserves injectively bounded dualizing complexes and ω_{π} is dualizing and since it is a line bundle, Bl is Gorenstein. The same proof holds for Z, since ι^{!}(O_{X}) is ω_{ι} (shifted) which is also a line bundle by Prop. A.11. For regularity, see [18, Thm. 1.19].

Proof of Theorem 2.6. Note that all the assumptions of Theorem 5.1 are fulfilled in the regular case, that is, in the setting of Section 2. Indeed, if X and Y are regular, Bl and U are regular, and the dualizing complexes on X, Y, Bl and U are simply shifted line bundles. The morphism α^{∗} : Pic(Y) → Pic(U) is then an isomorphism (homotopy invariance) and α̃ is automatically of finite Tor dimension, as any morphism to a regular scheme. Finally, the sequence on Picard groups is exact by Proposition A.3. Let K_{X} = L be the chosen line bundle on X. Then set L_{U} := K_{U} = υ^{∗}L and choose L_{Y} = K_{Y} to be the unique line bundle (up to isomorphism) such that α^{∗}L_{Y} = L_{U}. By (8), we have α̃^{∗}L_{Y} = π^{∗}L ⊗ O(E)^{⊗λ(L)} = π^{!}L ⊗ O(E)^{⊗(λ(L)−c+1)}, where the last equality holds since π^{!}L = O(E)^{⊗(c−1)} ⊗ π^{∗}L by Proposition A.11 (vi). In other words, we have proved that α̃^{∗}K_{Y} = π^{!}K_{X} ⊗ O(E)^{⊗(λ(L)−c+1)}. In Theorem 5.1, we can then take n = λ(L) − c + 1 and the parity condition becomes λ(L) ≡ c−1 mod 2 for Case (A) and λ(L) ≡ c mod 2 for Case (B). So, Case (A) is the trivial one and corresponds to Theorem 2.3. Case (B) exactly gives Theorem 2.6 up to the identifications of line bundles explained in Appendix A.

A Line bundles and dualizing complexes

We use Hartshorne [15] or Liu [18] as general references for algebraic geometry. We still denote by Sch the category of noetherian separated connected schemes (we do not need “over ℤ[\frac{1}{2}]” in this appendix).

Definition A.1. Let ι̃: E ↪ B be a regular closed immersion of codimension one, with B ∈ Sch. Consider the ideal I_{E} ⊂ O_{B} defining E

\xymatrix{0 \ar[r] & I_{E} \ar[r]^{σ_{E}} & O_{B} \ar[r] & ι̃_{∗}O_{E} \ar[r] & 0}.  (22)

By assumption, I_{E} is an invertible ideal, i.e. a line bundle. The line bundle associated to E is defined as its dual O(E) := (I_{E})^{∨}, see [15, II.6.18]. We thus have by construction a global section σ_{E} : O(E)^{∨} → O_{B}, which vanishes exactly on E. This gives an explicit trivialization of O(E) outside E. On the other hand, the restriction of O(E) to E is the normal bundle O(E)_{|E} ≅ N_{E/B}.

Example A.2. Let Bl = Bl_{Z}(X) be the blow-up of X along a regular closed immersion Z ↪ X as in Setup 1.1. Let I = I_{Z} ⊂ O_{X} be the sheaf of ideals defining Z. By construction of the blow-up, we have Bl = Proj(S) where S is the sheaf of graded O_{X}-algebras

S := O_{X} ⊕ I ⊕ I^{2} ⊕ I^{3} ⊕ ⋯

Similarly, E = Proj(S/J) where J := I · S ⊂ S is the sheaf of homogeneous ideals

J = I ⊕ I^{2} ⊕ I^{3} ⊕ I^{4} ⊕ ⋯

So, E = ℙ_{Z}(C_{Z/X}) is a projective bundle over Z associated to the vector bundle C_{Z/X} = I/I^{2} which is the conormal bundle of Z in X. Associating O_{Bl}-sheaves to graded S-modules, the obvious exact sequence 0 → J → S → S/J → 0 yields

\xymatrix{0 \ar[r] & J̃ \ar[r]^{σ_{E}} & O_{Bl} \ar[r] & ι̃_{∗}O_{E} \ar[r] & 0}.  (23)

Compare (22). This means that here I_{E} = J̃. But now, J is obviously S(1) truncated in non-negative degrees. Since two graded S-modules which coincide above some degree have the same associated sheaves, we have I_{E} = J̃ = S̃(1) = O_{Bl}(1). Consequently, O(E) = (I_{E})^{∨} = O_{Bl}(−1). In particular, we get

O(E)_{|E} = O_{Bl}(−1)_{|E} = O_{E}(−1).  (24)

Proposition A.3 (Picard group in codimension one). Let B ∈ Sch be a scheme and ι̃ : E ↪ B be a regular closed immersion of codimension one of an irreducible subscheme E ∈ Sch with open complement υ̃ : U ↪ B. We then have a complex

\xymatrix{ℤ \ar[r] & Pic(B) \ar[r]^{υ̃^{∗}} & Pic(U)}  (25)

where the first map sends 1 to the line bundle O(E) associated to E. This complex is exact if B is normal, and υ̃^{∗} is surjective when B is furthermore regular. It is also exact when B is the blow-up of a normal scheme X along a regular embedding.

Proof. (25) is a complex since O(E) is trivial on U. When B is normal, Pic(B) injects in the group Cl(B) of Weil divisors (see [18, 7.1.19 and 7.2.14 (c)]), for which the same sequence holds by [15, Prop. II.6.5]. Exactness of (25) then follows by diagram chase. The surjectivity of υ̃^{∗} when B is regular follows from [15, Prop. II.6.7 (c)]. When B is the blow-up of X along Z, we can assume that codim_{X}(Z) ≥ 2 by the previous point. Then, the result again follows by diagram chase, using that Pic(B) = Pic(X)⊕ℤ, as proved in Proposition A.6 (i) below.

Remark A.4. Note that the blow-up of a normal scheme along a regular closed embedding isn’t necessarily normal if the subscheme is not reduced. For example, take X = 𝔸^{2} = Spec(k[x, y]) and Z defined by the equations x^{2} = y^{2} = 0. Then, Bl is the subscheme of 𝔸^{2} × ℙ^{1} defined by the equations x^{2}v = y^{2}u where [u : v] are homogeneous coordinates for ℙ^{1} and it is easy to check that the whole exceptional fiber is singular. Thus Bl is not normal (not even regular in codimension one).

Proposition A.5 (Picard group of a projective bundle). Let X ∈ Sch be a (connected) scheme and F a vector bundle over X. We consider the projective bundle ℙ_{X}(F) associated to F. Its Picard group is Pic(X) ⊕ ℤ where ℤ is generated by O(−1) and Pic(X) comes from the pull-back from X.

Proof. Surjectivity of Pic(X) ⊕ ℤ → Pic(ℙ_{X}(F)) is a formal consequence of Quillen’s formula [23, Prop. 4.3] for the K-theory of a projective bundle. Indeed, the determinant map K_{0} → Pic is surjective with an obvious set theoretic section and can easily be computed on each component of Quillen’s formula. Injectivity is obtained by pulling back to the fiber of a point for the ℤ component, and by the projection formula for the remaining Pic(X) component.

Proposition A.6 (Picard group of a blow-up). Under Setup 1.1, we have :

(i) The Picard group of Bl = Bl_{Z}(X) is isomorphic to Pic(X) ⊕ ℤ where the direct summand Pic(X) comes from the pull-back π^{∗} and ℤ is generated by the class of the exceptional divisor O(E) = O_{Bl}(−1).

(ii) If X is normal, the map υ^{∗} : Pic(X) → Pic(U) is injective. If X is regular it is an isomorphism.

(iii) The exceptional fiber E is the projective bundle ℙ(C_{Z/X}) over Z and its Picard group is therefore Pic(Z) ⊕ ℤ where ℤ is generated by O_{E}(−1).

(iv) The pull-back ι̃^{∗} : Pic(Bl) → Pic(E) maps [O(E)] ∈ Pic(Bl) to [O_{E}(−1)].

Under these identifications, Diagram (1) induces the following pull-back maps on Picard groups :

\xymatrix{Pic(Z) \ar[d]_{_{( \begin{matrix}id \\ 0 \end{matrix} )}} & Pic(X) \ar[l]_{ι^{∗}} \ar[r]^{υ^{∗}} \ar[d]^{_{( \begin{matrix}id \\ 0 \end{matrix} )}} & Pic(U) . \\ Pic(Z) ⊕ ℤ & Pic(X) ⊕ ℤ \ar[l]_{_{( \begin{matrix}ι^{∗} & 0 \\ 0 & id \end{matrix} )}} \ar[ur]_{(υ^{∗} 0)}}

Proof. By Example A.2, we get (iv) and we can deduce (iii) from Proposition A.5. To prove (ii), use that for X normal (resp. regular) Pic(X) injects into (resp. is isomorphic to) the group Cl(X) of Weil divisors classes (see [18, 7.1.19 and 7.2.14 (c), resp. 7.2.16]), and that Cl(X) = Cl(U) since codim_{X}(Z) ≥ 2. Finally, for (i), consider the commutative diagram

\xymatrix{D_{perf}(Z) \ar[d]_{_{L_{π̃^{∗}}}} & D_{perf}(X) \ar[l]_{Lι^{∗}} \ar[d]^{Lπ^{∗}} \\ D_{perf}(E) & D_{perf}(Bl) \ar[l]_{_{L_{ι̃^{∗}}}}}  (26)

of induced functors on the derived categories of perfect complexes. We will use : Fact 1 : The tensor triangulated functors Lπ^{∗} and Lπ̃^{∗} are fully faithful with left inverse R π_{∗} and R π̃_{∗} respectively, see Thomason [24, Lemme 2.3]. Fact 2 : If M ∈ D_{perf}(Bl) is such that Lι̃^{∗}(M) ≃ Lπ̃^{∗}(N) for some N ∈ D_{perf}(Z), then M ≃ Lπ^{∗}(L) for some L ∈ D_{perf}(X), which must then be R π_{∗}(M) by Fact 1. This follows from [11, Prop. 1.5]. (In their notation, our assumption implies that M is zero in all successive quotients D_{perf}^{i+1}(Bl)/ D_{perf}^{i}(Bl) hence belongs to D_{perf}^{0}(Bl).)

Hence Pic(X) ⊕ ℤ → Pic(Bl) is injective : If L is a line bundle on X and n ∈ ℤ are such that Lπ^{∗}(L) ⊗ O_{Bl}(n) is trivial then we get n = 0 by restricting to E and applying (iii), and we get L ≃ R π_{∗} Lπ^{∗}L ≃ R π_{∗}O_{Bl} ≃ R π_{∗} Lπ^{∗}O_{X} ≃ O_{X} by Fact 1. So, let us check surjectivity of Pic(X) ⊕ ℤ → Pic(Bl). Let M be a line bundle on Bl. Using (iii) again and twisting with O_{Bl}(n) if necessary, we can assume that Lι̃^{∗}(M) is isomorphic to Lπ̃^{∗}N = π̃^{∗}N for some line bundle N on Z. By Fact 2, there exists L ∈ D_{perf}(X) such that Lπ^{∗}(L) ≃ M. It now suffices to check that this L ∈ D_{perf}(X) is a line bundle. The natural (evaluation) map L^{∨} ⊗ L → O_{X} is an isomorphism, since it is so after applying the fully faithful tensor functor Lπ^{∗} : D_{perf}(X) → D_{perf}(Bl). So L ∈ D_{perf}(X) is an invertible object, hence it is the m^{th} suspension of a line bundle for m ∈ ℤ, see [7, Prop. 6.4]. Using (26), one checks by restricting to Z that m = 0, i.e. L is a line bundle.

* * * We now discuss dualizing complexes and relative canonical bundles. First of all, we mention the essential uniqueness of dualizing complexes on a scheme.

Lemma A.7. Let X ∈ Sch be a scheme admitting a dualizing complex K_{X}. Then :

(i) For any line bundle L and any integer i, the complex K_{X} ⊗ L[i] is also a dualizing complex and any dualizing complex on X is of this form.

(ii) If K_{X} ⊗ L[i] ≃ K_{X} in the derived category of X, for some line bundle L and some integer i, then L ≃ O_{X} and i = 0.

In other words, the set of isomorphism classes of dualizing complexes on X is a principal homogeneous space under the action of Pic(X) ⊕ ℤ.

Proof. For (i), see [19, Lemma 3.9]. Let us prove (ii). We have the isomorphisms

O_{X} →^{∼} RHom(K_{X}, K_{X}) ≃ RHom(K_{X}, K_{X} ⊗ L[n])

≃ RHom(K_{X}, K_{X}) ⊗ L[n]←^{∼}L[n]

in the coherent derived category. The first and last ones hold by [19, Prop. 3.6]. We thus obtain an isomorphism O_{X} ≃ L[n] in the derived category of perfect complexes (it is a full subcategory of the coherent one). This forces n = 0 and the existence of an honest isomorphism of sheaves O_{X} ≃ L, see [7, Prop. 6.4] if necessary.

We now use the notion of local complete intersection (l.c.i.) morphism, that is, a morphism which is locally a regular embedding followed by a smooth morphism, see [18, § 6.3.2]. The advantage of such morphisms f : X′ → X is that f^{!} is just Lf^{∗} twisted by a line bundle ω_{f} and shifted by the relative dimension dim(f).

Proposition A.8. Let f : X′ → X be an l.c.i. morphism with X, X′ ∈ Sch. Assume that f is proper. Then f^{!}(O_{X}) is a shifted line bundle ω_{f}[dim(f)] and there exists a natural isomorphism f^{!}(O_{X}) ⊗ Lf^{∗}(−) →^{∼} f^{!}(−). In particular, f^{!} preserves the subcategory D_{perf} of D_{coh}^{b}.

Proof. There is always a natural morphism f^{!}(O_{X}) ⊗ Lf^{∗}(−) → f^{!}(−). One shows that it is an isomorphism and that f^{!}(O_{X}) is a line bundle directly from the definition, since both these facts can be checked locally, are stable by composition and are true for (closed) regular immersions and smooth morphisms by Hartshorne [14, Ch. III]. The subcategory D_{perf} is then preserved since both Lf^{∗} and tensoring by a line bundle preserve it.

The above proposition reduces the description of f^{!} to that of the line bundle ω_{f}.

Proposition A.9. In the following cases, we have concrete descriptions of ω_{f}.

(i) When f : X′ → X is smooth and proper, ω_{f} ≃ det(Ω_{X′/X}^{1}) is the determinant of the sheaf of differentials. In particular, when f is the projection of a projective bundle ℙ(F) to its base, where F is a vector bundle of rank r, then ω_{f} ≃ f^{∗}(detF) ⊗ O_{ℙ(F)}(−r).

(ii) When f : X′ ↪ X is a regular closed immersion, ω_{f} ≃ det(N_{X′/X}) is the determinant of the normal bundle. In particular when f : E ↪ B is the inclusion of a prime divisor (Def. A.1), we have ω_{f} ≃ O(E)_{|E}.

Proof. See [25, Prop. 1 and Thm. 3]. See alternatively [18, § 6.4.2].

Remark A.10. All morphisms along which we consider push-forward in this article are l.c.i. It might not be obvious for π : Bl → X but this follows from [1, VII 1.8 p. 424] (it is locally of the form mentioned there). So, ω_{π} is also a line bundle. Let us now describe the relative canonical line bundles in terms of ω_{ι} = det(N_{Z/X}).

Proposition A.11. With the notation of Setup 1.1, we have

(i) ωι̃ = O(E)_{|E} = O_{E}(−1)

(ii) ωπ̃ = π̃^{∗}ω_{ι}^{∨} ⊗ O(E)_{|E}^{⊗c} = π̃^{∗}ω_{ι}^{∨} ⊗ O_{E}(−c)

(iii) ω_{π} = O(E)^{⊗(c−1)} = O_{Bl}(1 − c).

By Proposition A.8, it implies that we have

(iv) ι̃^{!}(−) = O(E)_{|E} ⊗ Lι̃^{∗}(−)[−1] = O_{E}(−1) ⊗ Lι̃^{∗}(−)[−1]

(v) π̃^{!}(−) = π̃^{∗}ω_{ι}^{∨}⊗O(E)_{|E}^{⊗c}⊗Lπ̃^{∗}(−)[c−1] = π̃^{∗}ω_{ι}^{∨}⊗O_{E}(−c)⊗Lπ̃^{∗}(−)[c−1]

(vi) π^{!}(−) = O(E)^{⊗(c−1)} ⊗ Lπ^{∗}(−) = O_{Bl}(1 − c) ⊗ Lπ^{∗}(−).

Proof. Points (i) and (ii) follow from Proposition A.9 (ii) and (i), respectively. They imply (iv) and (v). To prove point (iii) let us first observe that the exact sequence (22) gives rise to an exact triangle

O_{Bl}(l + 1) → O_{Bl}(l) → R ι̃_{∗}(O_{E}(l)) → O_{Bl}(l + 1)[1]

in D_{perf}(Bl) for any l ∈ ℤ. Applying R π_{∗} to this triangle and using that

R π_{∗} R ι̃_{∗}O_{E}(l) = R ι_{∗} R π̃_{∗}O_{E}(l) = 0 for − c < l < 0

(by [13, 2.1.15]), we obtain by induction that R π_{∗}O_{Bl}(l) = R π_{∗}O_{Bl} = O_{X} for −c < l ≤ 0. In particular R π_{∗}O_{Bl}(1 − c) = O_{X}. We now use the filtration of [11, Prop. 1.5]. Let us show that π^{!}(O_{X}) ⊗ O_{Bl}(c − 1) is in D_{perf}^{0}(Bl). By loc. cit. it suffices to show that Lι̃^{∗}(π^{!}(O_{X}) ⊗ O_{Bl}(c − 1)) is in D_{perf}^{0}(E). It follows from the sequence of isomorphisms

Lι̃^{∗}(π^{!}(O_{X}) ⊗ O_{Bl}(c − 1)) ≃ Lι̃^{∗}π^{!}(O_{X}) ⊗ O_{E}(c − 1) ≃^{(iv)} ι̃^{!}π^{!}(O_{X}) ⊗ O_{E}(c)[1]

≃ π̃^{!}ι^{!}(O_{X}) ⊗ O_{E}(c)[1] ≃^{A.8} π̃^{!}(ω_{ι}) ⊗ O_{E}(c)[−c + 1] ≃^{(v)} O_{E} .

Since L := π^{!}(O_{X}) ⊗ O_{Bl}(c − 1) is in D_{perf}^{0}(Bl), it is of the form Lπ^{∗}M for M = R π_{∗}L (by Fact 1 in the proof of Prop. A.6) which we compute by duality:

R π_{∗}(π^{!}(O_{X}) ⊗ O_{Bl}(c − 1)) ≃ R π_{∗}RHom(O_{Bl}, π^{!}(O_{X}) ⊗ O_{Bl}(c − 1)) ≃

≃ R π_{∗}RHom(O_{Bl}(1 − c), π^{!}(O_{X})) ≃^{(†)} RHom(R π_{∗}O_{Bl}(1 − c), O_{X}) ≃^{(⋆)}

≃ RHom(O_{X}, O_{X}) ≃ O_{X}

where (⋆) is by the computation at the beginning of the proof, (†) is the duality isomorphism and all other isomorphisms are obtained as consequences of the monoidal structure on the D_{coh}^{b} involved (see [10] and [8] for details). Hence, π^{!}(O_{X}) ≃ O_{Bl}(1 − c) as announced. This proves (iii) and thus (vi).

Finally, we also use dualizing complexes in the context of an 𝔸^{∗}-bundle U → Y, i.e. a morphism that is locally of the form 𝔸_{Y}^{n} → Y (and is in particular flat).

Lemma A.12. Let α : U → Y be an 𝔸^{∗}-bundle. Assume that Y admits a dualizing complex K_{Y}. Then Lα^{∗}(K_{Y}) = α^{∗}(K_{Y}) is a dualizing complex on U.

Proof. This can be checked locally, so we can assume α decomposes as α = f ∘u for an open immersion u : 𝔸_{Y}^{n} ↪ ℙ_{Y}^{n} followed by the structural projection f : ℙ_{Y}^{n} → Y. Note that α, u and f are all flat. We have by Propositions A.8 and A.9 (i)

u^{∗}f^{!}K_{Y} [n] = u^{∗}(O(−n − 1) ⊗ f^{∗}K_{Y}) ≃ u^{∗}f^{∗}K_{Y} = α^{∗}K_{Y}

where the second equality comes from the triviality of O(−n − 1) on 𝔸_{Y}^{n}. Now u^{∗}f^{!}K_{Y} [n] is dualizing because proper morphisms, open immersions and shifting preserve dualizing complexes.

Acknowledgments : We thank Marc Levine for precious discussions on Theorem 1.3 and Example 1.5, Jean Fasel for a florilegium of relative bundles and Burt Totaro for several discussions on algebraic geometry.

References

[1] Théorie des intersections et théorème de Riemann-Roch, Séminaire de Géométrie Algébrique du Bois-Marie 1966–1967 (SGA 6), dirigé par P. Berthelot, A. Grothendieck et L. Illusie. Lecture Notes in Mathematics, Vol. 225. Springer-Verlag, Berlin, 1971.

[2] J. Arason, Der Wittring projektiver Räume, Math. Ann. 253 (1980), no. 3, 205–212.

[3] P. Balmer, Triangular Witt Groups Part 1: The 12-Term Localization Exact Sequence, K-Theory 4 (2000), no. 19, 311–363.

[4] , Products of degenerate quadratic forms, Compos. Math. 141 (2005), no. 6, 1374–1404.

[5] , Witt groups, Handbook of K-theory. Vol. 2, Springer, Berlin, 2005, pp. 539–576.

[6] P. Balmer and B. Calmès, Witt groups of Grassmann varieties, preprint, 2008.

[7] P. Balmer and G. Favi, Gluing techniques in triangular geometry, Q J Math 58 (2007), no. 4, 415–441.

[8] B. Calmès and J. Hornbostel, Push-forwards for Witt groups of schemes, arXiv:0806.0571, 2008, to appear in Comment. Math. Helv.

[9] , Witt motives, transfers and dévissage, preprint, available at

http://www.math.uiuc.edu/K-theory/0786/, 2006.

[10] B. Calmes and J. Hornbostel, Tensor-triangulated categories and dualities, Theory Appl. Categ. 22 (2009), No. 6, 136–198.

[11] G. Cortiñas, C. Haesemeyer, M. Schlichting, and C. Weibel, Cyclic homology, cdh-cohomology and negative K-theory, Ann. of Math. 167 (2008), 549–573.

[12] S. Gille, Homotopy invariance of coherent Witt groups, Math. Z. 244 (2003), no. 2, 211–233.

[13] A. Grothendieck, Éléments de géométrie algébrique. III. Étude cohomologique des faisceaux cohérents. I, Inst. Hautes Études Sci. Publ. Math. (1961), no. 11, 167.

[14] R. Hartshorne, Residues and duality, Lecture notes of a seminar on the work of A. Grothendieck, given at Harvard 1963/64. With an appendix by P. Deligne. Lecture Notes in Mathematics, No. 20, Springer-Verlag, Berlin, 1966.

[15] , Algebraic geometry, Springer-Verlag, New York, 1977, Graduate Texts in Mathematics, No. 52.

[16] D. Laksov, Algebraic Cycles on Grassmann Varieties, Adv. in Math. 9 (1972), no. 3, 267–295.

[17] M. Levine and F. Morel, Algebraic cobordism, Springer Monographs in Mathematics, Springer, Berlin, 2007.

[18] Q. Liu, Algebraic geometry and arithmetic curves, Oxford Graduate Texts in Mathematics, vol. 6, Oxford University Press, Oxford, 2002.

[19] A. Neeman, Derived categories and Grothendieck duality, preprint 791 of

http://www.crm.es/Publications/Preprints08.htm.

[20] A. Nenashev, Projective push-forwards in the Witt theory of algebraic varieties, Adv. Math. 220 (2009), no. 6, 1923–1944.

[21] A. Nenashev and K. Zainoulline, Oriented cohomology and motivic decompositions of relative cellular spaces, J. Pure Appl. Algebra 205 (2006), no. 2, 323–340.

[22] I. Panin, Riemann-Roch theorems for oriented cohomology, Axiomatic, enriched and motivic homotopy theory, NATO Sci. Ser. II Math. Phys. Chem., vol. 131, Kluwer Acad. Publ., Dordrecht, 2004, pp. 261–333.

[23] D. Quillen, Higher algebraic K-theory: I, Algebraic K-theory, Lecture Notes in Math., no. 341, Springer-Verlag, 1973, pp. 83–147.

[24] R. W. Thomason, Les K-groupes d’un schéma éclaté et une formule d’intersection excédentaire, Invent. Math. 112 (1993), no. 1, 195–215.

[25] J.-L. Verdier, Base change for twisted inverse image of coherent sheaves, Algebraic Geometry (Internat. Colloq., TIFR, Bombay, 1968), Oxford Univ. Press, London, 1969, pp. 393–408.

[26] C. Walter, Grothendieck-Witt groups of projective bundles, preprint, 2003.

Paul Balmer Department of Mathematics UCLA Los Angeles, CA 90095-1555 USA balmer@math.ucla.edu Baptiste Calmès LML Fac. des Sciences J. Perrin Université d’Artois 62307 Lens France calmes@math.jussieu.fr